OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 5x - 3y = -11 2x - 6y = -14

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

Can you try to explain your work more? I got lost when I was trying to do this one.

OpenStudy (trojanpoem):

Ok, take the 2 equations: to eliminate one of the variables , they must have the same coefficient , so we will multiply each equation times the coefficient of the variable we want to eliminate -30 x + 18y = 66 6x - 18 y = -42 -24x = 24 x = -1 now plug x in any simple equation 5x - 3y = -11 -5 -3y = -11 -3y = -6 y = 2

OpenStudy (anonymous):

Thank you so much!! I think I get it now! Do you know how to figure out equations with 3 variables?

OpenStudy (denisaboichuk):

y=2